30,501.39J of heat is added to 1000 g of ice at -10 oC.  How much ice melts?
Answer #1

To bring it to 0 degree = mct = 1*2108*10=21080J

Heat Left= 9421J

Now 9421=mL

L=333J/g

therefore Mass= 9421/333 = 28.3 gms (Approx)